sheri's age in 20 years will be the same as terry's age is now. then ten years from now, terry's age will be twice as sheri's. what are their present ages?
Answer by Theo(13342)

x = sheri's age now.
y = terry's age now.

terry's age now is the same as what sheri's age will be 20 years from now, therefore:

y = x + 20

in 10 years, terry's age will be twice sheri's age will be in 10 years.

y + 10 = 2 * (x + 10)

since y = x + 20, replace y with x + 20 and that equation becomes:

x + 20 + 10 = 2 * (x + 10)

combine like terms and simplify to get x + 30 = 2 * x + 20

subtract x from both sides of this equation and subtract 20 from both sides of this equation to get 10 = x.

that's sheri's age now.

since y = x + 20, then y = 30.

that's terry's age now.

you have:

sheri is 10 years old now.
terry is 30 years old now.

in 20 years sherie will be the same age as terry is now.
since 10 + 20 = 30, this is a true statement.

in 10 years, terry will be 40 years old.
in 10 years, sheri will be 20 years old.
40 is twice as old as 20, therefore the statement that terry will be twice as old as sheri is also a true statement.

the solution is that sheri is 10 years old now and terry is 30 years old now.